# Learning Primitive Embodied World Models: Towards Scalable Robotic Learning

### v1

@@ 1 Introduction
Embodied agents capable of planning and decision-making in complex environments rely heavily on internal representations of the world, commonly referred to as world models. Recent advances in pretrained video generation models–powered by self-supervised learning on vast internet-scale datasets–have demonstrated remarkable capabilities in synthesizing fine-grained, temporally coherent visual sequences from natural language descriptions Du et al. (2023).
These successes have inspired a growing body of work to leverage such models as world models for embodied agents, casting policy learning as a video generation problem Ko et al. (2023); Du et al. (2023); Yang et al. (2023); Brooks et al. (2024); Black et al. (2024a); Xiang et al. (2024); Bruce et al. (2024); Hu et al. (2024); Zhou et al. (2024); Dalal et al. (2025); Liang et al. (2024a); Clark et al. (2025); Hu et al. (2024); Li et al. (2025); Team et al. (2025a); Pertsch et al. (2025); Zhen et al. (2025); Zhu et al. (2025); Hafner et al. (2025): given a language-specified goal, the model generates a future video roll-out, from which actions are subsequently extracted. This paradigm offers compelling advantages in interpretability, generalization, and unified representation across diverse environments.
@@
Figure 1: While densely distributed data can enable broad generalization, embodied data often suffer from sparsity Du and Kaelbling (2024); Xue et al. (2025b). The rightmost schematic highlights how organizing embodied data at the primitive level–along orthogonal dimensions such as action and object–supports compositional generalization even under limited data availability.
@@ 2.1 The Motivation and Definition of Primitive Embodied Data
Collecting real-world embodied data is labor-intensive. Existing datasets, despite size, suffer from limited generalization beyond minor scene variations (Xue et al., 2025b; Brohan et al., 2022; Collaboration et al., 2023; Khazatsky et al., 2024b; Walke et al., 2023; Dalal et al., 2025). To improve generalization in embodied agents, we advocate organizing and collecting data at the level of semantically grounded primitives:
@@ 2.4 The High Potential of Leveraging Action-Free Video Data for Scalable, Web-Scale Embodied Learning
It is worth noting that the LIBERO Liu et al. (2024a) data we use not only includes the original dataset’s multi-view replays but also videos generated through OpenVLA rollouts as training data. This directly demonstrates our method’s ability to leverage any embodied video–whether from rollouts of other policies or from web-scale data–to improve model performance, highlighting one of the key advantages of using video as a universal representation.
@@ 3.1 Dual-Level Compositional Generalization
Implicit Compositionality of Diffusion Video Generation Diffusion models are capable of generating photo-realistic images that combine elements which likely do not appear together in the training set, demonstrating the ability to compositionally generalize Liu et al. (2022a); Liang et al. (2024b).
@@ 3.2 Training Strategy
Sim-Real Hybrid Data Strategy In practice, we found that when training solely on real-world data, the model struggles to capture fine details of fast-moving parts such as the gripper, due to the reconstruction loss in the VAE. To enable the model to learn cleaner proprioceptive motion patterns, we introduced simulation data, including data from RLBench James et al. (2020a) and LIBERO Liu et al. (2024a), and adopt a sim-real hybrid data mixing strategy. For detailed data information, please refer to Appendix D.1. This approach proves highly effective: the final model successfully combines the rich, complex textures from real-world data with the precise and clear kinematic motions from simulation, significantly enhancing the overall generation quality. Appendix I.3 Figure 12 shows a comprehensive comparison.
@@ 3.3 Causal Distillation and Acceleration: Real-Time On-the-Fly Future Frame Prediction
To address this, we adopt causal video generation via knowledge distillation, following CausVid Yin et al. (2025). We train a student model to predict future frames in a strictly causal manner–each frame is generated based on past observations only, without access to future context. The student performs only 4 denoising steps per chunk (with 4 frames), enabling low-latency inference. To maintain generation quality under aggressive acceleration, we employ self-forcing–using the model’s own past predictions as input for future steps, closing the loop between prediction and conditioning.
@@ 4.1 Application 1: High-Quality Embodied Video Generation Facilitating Direct 6-DoF Trajectory Extraction
We apply Gen6D Liu et al. (2022b), an off-the-shelf, zero-shot 6-DoF pose estimator, to the generated video frames to recover the full 6-DoF motion trajectory of the robot end-effector. This enables direct mapping from visual prediction to actionable control signals–bridging the gap between generative world models and real robot execution. Technical details of this extraction pipeline are provided in Appendix B. The experimental results are provided in Appendix B.2.
@@ 4.2 Application 2: Plug-and-Play Long-Horizon Compositional Generalization
Table 1: Overall success rate on RLBench tasks. We compare our method against existing baselines on 9 manipulation tasks from the RLBench benchmark James et al. (2020b), with success rates averaged over 100 episodes. Results for Image-BC (Jang et al., 2022), UniPi (Du et al., 2023), and 4DWM Zhen et al. (2025) are directly cited from the 4DWM paper Zhen et al. (2025). Our method (bottom row) achieves the highest success rate on most tasks, with consistent gains in articulated and contact-rich scenarios.
@@ 5.1 Performance
Planning Our method performs a notable success rate, attributable to the effective decomposition of tasks into manageable primitives. The test scenarios outlined in Table 2 were not present in the training set, yet our method achieves a high primitive accuracy of 18 out of 20 for the task of picking up a cup. Similar performance is observed across other tasks, with accuracy rates of 16 out of 20 for moving cloth and 15 out of 20 for folding cloth. This robust planning capability ensures that the robot can generate reliable sequences of actions, setting a strong foundation for successful execution. Notably, OpenVLA fails completely in the zero-shot setting, highlighting its reliance on task-specific adaptation. In contrast, our method achieves strong generalization without retraining the video model, benefiting from modularity and spatial grounding in its design.
@@ 6 Related Work
Video Generation as World Models Given the overwhelming success of Diffusion Models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021, 2022) in image generation (Rombach et al., 2022b; Ramesh et al., 2022; Ding et al., 2022), video diffusion models (Li et al., 2022; Ho et al., 2022; He et al., 2023; Li et al., 2023a; Blattmann et al., 2023; Singer et al., 2022; Wang et al., 2023; Chen et al., 2023, 2024a; Xing et al., 2023) have emerged as a prominent research focus. With the emergence of Sora (Brooks et al., 2024; Liu et al., 2024e), Transformer-based video generation has gained traction as a mainstream paradigm. Leveraging the Diffusion Transformer (DiT) architecture Peebles and Xie (2023), these models significantly improve video stability and temporal consistency (Zheng et al., 2024; Lin et al., 2024; Yang et al., 2024c; Kong et al., 2024; Polyak et al., 2025; Yang et al., 2025b; Wang et al., 2025). Empowered by strong visual priors, video diffusion models show promising potential as pixel-level world models. A world model learns to predict future system states, either in a low-dimensional latent space (Achille and Soatto, 2017; Ferns et al., 2004; Lesort et al., 2018) or directly in pixel space (Chiappa et al., 2017; Ha and Schmidhuber, 2018; Hafner et al., 2020; Micheli et al., 2022; Chen et al., 2022). A large body of work (Ko et al., 2023; Du et al., 2023; Yang et al., 2023; Brooks et al., 2024; Black et al., 2024a; Xiang et al., 2024; Bruce et al., 2024; Hu et al., 2024; Zhou et al., 2024; Dalal et al., 2025; Liang et al., 2024a; Clark et al., 2025; Hu et al., 2024; Li et al., 2025; Team et al., 2025a; Pertsch et al., 2025; Zhen et al., 2025; Zhu et al., 2025; Hafner et al., 2025) has framed video generation as a new methodology for learning world models. Latest works, typically DiT-based and following a unified generation paradigm (Li et al., 2025; Team et al., 2025a; Zhen et al., 2025), are costly to train and lack real-time responsiveness. While hierarchical approaches (Ajay et al., 2023; Dalal et al., 2025; Zhou et al., 2024) offer structure, they struggle with real-time adaptation in dynamic environments and rigid cross-modal alignment. Error accumulation in layered designs remains an open problem. Several approaches (Ajay et al., 2023; Dalal et al., 2025; Zhou et al., 2024) adopt hierarchical architectures, among which Ajay et al. (2023); Dalal et al. (2025) are the closest to ours. However, these methods lack the integration of vision-language reasoning, planning, and grounding as in our approach. In contrast, our approach is distinguished by its ability to perform zero-shot policy execution, without relying on task-specific trained policies that often suffer from limited generalizability. Furthermore, our start-to-end point guidance surpasses their interactive refinement or reliance on 3D models in both efficiency and explainability. Some additional studies Seo et al. (2023); Wu et al. (2024); Yang et al. (2024b); Cheang et al. (2024) investigated non-diffusion approaches to modeling embodied world models.
@@
End-to-End Vision-Language-Action Learning End-to-end Vision-Language-Action (VLA) models constitute another major branch of embodied learning, directly mapping visual and linguistic inputs to actions without explicit dynamics modeling. Recent advances in vision-language models (VLMs) Chen et al. (2024c); Li et al. (2024a); Beyer et al. (2024); Li et al. (2024b) have catalyzed the development of unified VLA systems that integrate perception, grounding, reasoning, and control within a single architecture–departing from traditional modular pipelines. At the core of most VLA systems lies the intuition that pre-trained vision-language models (VLMs) can serve as high-level semantic engines. Methods like RT-2 Brohan et al. (2023b), OpenVLA Kim et al. (2024a), and RoboAgent Bharadhwaj et al. (2023) inject Internet-scale knowledge into robot policies by fine-tuning VLMs or conditioning them with low-level control tokens. Based on their action modeling paradigms, VLA approaches can be broadly categorized into regression-based Li et al. (2023b); Brohan et al. (2023a); Liu et al. (2024c), generative/diffusion-based Black et al. (2024b); Liu et al. (2024d); Team et al. (2024), and hybrid methods Liu et al. (2025a); Ye et al. (2024). While these unified policies show strong performance on short-horizon or single-task scenarios, they often struggle with generalization and scalability–particularly across long horizons, domains, or embodiments. To address these limitations, recent efforts have introduced hierarchical structures into end-to-end VLA models, either by modularizing perception and policy Liu et al. (2024b); Team et al. (2025b); Bjorck et al. (2025), or by structuring execution into multi-stage skills Shi et al. (2025); Physical Intelligence (2025); Raj et al. (2024). Though often labeled hierarchical, many still preserve end-to-end differentiability or latent consistency across stages. Such designs improve sample efficiency, handle long-horizon dependencies, and provide interpretable intermediate reasoning. Despite these advances, existing VLA systems remain heavily dependent on in-domain demonstrations. Some mitigate this via cross-modal or cross-embodiment data Intelligence et al. (2025); NVIDIA et al. (2025), or by scaling human video demonstrations Yang et al. (2024a); Tian et al. (2024), yet diversity and coverage remain limited. Our work enables zero-shot generalization across unseen tasks and settings, while maintaining modularity, interpretability, and execution efficiency.
@@ 7.1 Brief Summary of Methodology Advantages
Against Vanilla Video Diffusion Models-Based Methods Compared to long-horizon video diffusion models, our method operates at the primitive level with short-horizon prediction, drastically reducing computational cost and optimization difficulty. Leveraging causal distillation and self-forcing, we achieve real-time generation (12 FPS) in just 4 denoising steps, enabling online, on-the-fly rollout. This causal, modular design avoids error accumulation and supports dynamic, adaptive control–critical for real-world robotic deployment.
@@ B.1 The Gen6D Method
Specifically, given a generated video sequence $V=\{I_{1},I_{2},\ldots,I_{T}\}$ and a reference video $V_{r}=\{I_{r1},I_{r2},\ldots,I_{rT^{\prime}}\}$ of the robotic gripper, we estimate the gripper’s 6-DoF pose using a model-free, RGB-based pose estimator: $\mathbf{R},\mathbf{T}=PE(V,V_{r})$, where $\mathbf{R}$ and $\mathbf{T}$ denote the rotation and translation matrices, respectively. In our implementation, we use Gen6D [Liu et al., 2022b] for its simplicity and generalization capability.
@@
After generating the future rollout $x_{1:T}^{\text{img}}$ for a primitive $a_{k}$, we extract a 6-DoF end-effector trajectory $\tau_{k}=\{p_{1},\ldots,p_{T}\}$ using 6-DoF pose estimation Liu et al. [2022b] and geometric transformation. These trajectories are then mapped into Cartesian control space and executed by the robot.
@@ Appendix C Details for Hierachical Plug-and-Play Long-Horizon Compositional Generalization
We leverage LoRA-based adaptation to enable reasoning and spatial grounding, following a Chain-of-LoRA [Liu et al., 2025b]-style architecture. The structure is detailed in Figure 8.

### v2

@@ 1 Introduction
Figure 1: While densely distributed data can enable broad generalization, embodied data often suffer from sparsity (Du & Kaelbling, 2024; Xue et al., 2025b). The rightmost schematic highlights how organizing embodied data at the primitive level–along orthogonal dimensions such as action and object–supports compositional generalization even under limited data availability.
Embodied agents capable of planning and decision-making in complex environments rely heavily on internal representations of the world, commonly referred to as world models. Recent advances in pretrained video generation models–powered by self-supervised learning on vast internet-scale datasets–have demonstrated remarkable capabilities in synthesizing fine-grained, temporally coherent visual sequences from natural language descriptions (Du et al., 2023).
These successes have inspired a growing body of work to leverage such models as world models for embodied agents, casting policy learning as a video generation problem (Ko et al., 2023; Du et al., 2023; Yang et al., 2023; Brooks et al., 2024; Black et al., 2024a; Xiang et al., 2024; Bruce et al., 2024; Hu et al., 2024; Zhou et al., 2024; Dalal et al., 2025; Liang et al., 2024a; Clark et al., 2025; Hu et al., 2024; Li et al., 2025; Team et al., 2025a; Pertsch et al., 2025; Zhen et al., 2025b; c; a; Zhu et al., 2025; Hafner et al., 2025; Sun et al., 2024b; a; Ye et al., 2024a): given a language-specified goal, the model generates a future video roll-out, from which actions are subsequently extracted. This paradigm offers compelling advantages in interpretability, generalization, and unified representation across diverse environments.
@@ 2.1 The Motivation and Definition of Primitive Embodied Data
Collecting real-world embodied data is labor-intensive. Existing datasets, despite size, suffer from limited generalization beyond minor scene variations (Xue et al., 2025b; Brohan et al., 2022; Collaboration et al., 2023; Khazatsky et al., 2024b; Walke et al., 2023; Dalal et al., 2025).
To improve generalization in embodied agents, we advocate organizing and collecting data at the level of semantically grounded primitives:
@@ 2.4 The High Potential of Leveraging Action-Free Video Data for Scalable, Web-Scale Embodied Learning
It is worth noting that the LIBERO (Liu et al., 2024a) data we use not only includes the original dataset’s multi-view replays but also videos generated through OpenVLA rollouts as training data. This directly demonstrates our method’s ability to leverage any embodied video–whether from rollouts of other policies or from web-scale data–to improve model performance, highlighting one of the key advantages of using video as a universal representation.
@@ 3.1 Dual-Level Compositional Generalization
Implicit Compositionality of Diffusion Video Generation Diffusion models are capable of generating photo-realistic images that combine elements which likely do not appear together in the training set, demonstrating the ability to compositionally generalize (Liu et al., 2022a; Liang et al., 2024b).
@@ 3.2 Training Strategy
Sim-Real Hybrid Data Strategy In practice, we found that when training solely on real-world data, the model struggles to capture fine details of fast-moving parts such as the gripper, due to the reconstruction loss in the VAE. To enable the model to learn cleaner proprioceptive motion patterns, we introduced simulation data, including data from RLBench (James et al., 2020a) and LIBERO (Liu et al., 2024a), and adopt a sim-real hybrid data mixing strategy. For detailed data information, please refer to Appendix D.1.
This approach proves highly effective: the final model successfully combines the rich, complex textures from real-world data with the precise and clear kinematic motions from simulation, significantly enhancing the overall generation quality. Appendix I.3 Figure 12 shows a comprehensive comparison.
@@ 3.3 Causal Distillation and Acceleration: Real-Time On-the-Fly Future Frame Prediction
To address this, we adopt causal video generation via knowledge distillation, following CausVid (Yin et al., 2025). We train a student model to predict future frames in a strictly causal manner–each frame is generated based on past observations only, without access to future context. The student performs only 4 denoising steps per chunk (with 4 frames), enabling low-latency inference. To maintain generation quality under aggressive acceleration, we employ self-forcing–using the model’s own past predictions as input for future steps, closing the loop between prediction and conditioning.
@@ 4.1 Application 1: High-Quality Embodied Video Generation Facilitating Direct 6-DoF Trajectory Extraction
We apply Gen6D (Liu et al., 2022b), an off-the-shelf, zero-shot 6-DoF pose estimator, to the generated video frames to recover the full 6-DoF motion trajectory of the robot end-effector. This enables direct mapping from visual prediction to actionable control signals–bridging the gap between generative world models and real robot execution. Technical details of this extraction pipeline are provided in Appendix B. The experimental results are provided in Appendix B.2.
@@ 4.2 Application 2: Plug-and-Play Long-Horizon Compositional Generalization
Table 1: Overall success rate on RLBench tasks. We compare our method against existing baselines on 9 manipulation tasks from the RLBench benchmark (James et al., 2020b), with success rates averaged over 100 episodes. Results for Image-BC (Jang et al., 2022), UniPi (Du et al., 2023), and 4DWM (Zhen et al., 2025a) are directly cited from the 4DWM paper. Our method (bottom row) achieves the highest success rate on most tasks, with consistent gains in articulated and contact-rich scenarios.
@@ 5.1 Performance
Planning Our method performs a notable success rate, attributable to the effective decomposition of tasks into manageable primitives.
The test scenarios outlined in Table 2 were not present in the training set, yet our method achieves a high primitive accuracy of 18 out of 20 for the task of picking up a cup. Similar performance is observed across other tasks, with accuracy rates of 16 out of 20 for moving cloth and 15 out of 20 for folding cloth. This robust planning capability ensures that the robot can generate reliable sequences of actions, setting a strong foundation for successful execution. Notably, OpenVLA fails completely in the zero-shot setting, highlighting its reliance on task-specific adaptation. In contrast, our method achieves strong generalization without retraining the video model, benefiting from modularity and spatial grounding in its design.
@@ 6 Related Work
Video Generation as World Models Given the overwhelming success of Diffusion Models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021; 2022) in image generation (Rombach et al., 2022b; Ramesh et al., 2022; Ding et al., 2022), video diffusion models (Li et al., 2022; Ho et al., 2022; He et al., 2023; Li et al., 2023a; Blattmann et al., 2023; Singer et al., 2022; Wang et al., 2023; Chen et al., 2023; 2024a; Xing et al., 2023) have emerged as a prominent research focus. With the emergence of Sora (Brooks et al., 2024; Liu et al., 2024e), Transformer-based video generation has gained traction as a mainstream paradigm. Leveraging the Diffusion Transformer (DiT) architecture Peebles & Xie (2023), these models significantly improve video stability and temporal consistency (Zheng et al., 2024; Lin et al., 2024; Yang et al., 2024c; Kong et al., 2024; Polyak et al., 2025; Yang et al., 2025b; Wang et al., 2025). Empowered by strong visual priors, video diffusion models show promising potential as pixel-level world models. A world model learns to predict future system states, either in a low-dimensional latent space (Achille & Soatto, 2017; Ferns et al., 2004; Lesort et al., 2018) or directly in pixel space (Chiappa et al., 2017; Ha & Schmidhuber, 2018; Hafner et al., 2020; Micheli et al., 2022; Chen et al., 2022). A large body of work (Ko et al., 2023; Du et al., 2023; Yang et al., 2023; Brooks et al., 2024; Black et al., 2024a; Xiang et al., 2024; Bruce et al., 2024; Hu et al., 2024; Zhou et al., 2024; Dalal et al., 2025; Liang et al., 2024a; Clark et al., 2025; Hu et al., 2024; Li et al., 2025; Team et al., 2025a; Pertsch et al., 2025; Zhen et al., 2025a; Zhu et al., 2025; Hafner et al., 2025) has framed video generation as a new methodology for learning world models. Latest works, typically DiT-based and following a unified generation paradigm (Li et al., 2025; Team et al., 2025a; Zhen et al., 2025a), are costly to train and lack real-time responsiveness. While hierarchical approaches (Ajay et al., 2023; Dalal et al., 2025; Zhou et al., 2024) offer structure, they struggle with real-time adaptation in dynamic environments and rigid cross-modal alignment. Error accumulation in layered designs remains an open problem. Several approaches (Ajay et al., 2023; Dalal et al., 2025; Zhou et al., 2024) adopt hierarchical architectures, among which Ajay et al. (2023); Dalal et al. (2025) are the closest to ours. However, these methods lack the integration of vision-language reasoning, planning, and grounding as in our approach. In contrast, our approach is distinguished by its ability to perform zero-shot policy execution, without relying on task-specific trained policies that often suffer from limited generalizability. Furthermore, our start-to-end point guidance surpasses their interactive refinement or reliance on 3D models in both efficiency and explainability. Some additional studies Seo et al. (2023); Wu et al. (2024); Yang et al. (2024b); Cheang et al. (2024) investigated non-diffusion approaches to modeling embodied world models.
@@
End-to-End Vision-Language-Action Learning End-to-end Vision-Language-Action (VLA) models constitute another major branch of embodied learning, directly mapping visual and linguistic inputs to actions without explicit dynamics modeling. Recent advances in vision-language models (VLMs) Chen et al. (2024c); Li et al. (2024a); Beyer et al. (2024); Li et al. (2024b) have catalyzed the development of unified VLA systems that integrate perception, grounding, reasoning, and control within a single architecture–departing from traditional modular pipelines. At the core of most VLA systems lies the intuition that pre-trained vision-language models (VLMs) can serve as high-level semantic engines. Methods like RT-2 Brohan et al. (2023b), OpenVLA Kim et al. (2024a), and RoboAgent Bharadhwaj et al. (2023) inject Internet-scale knowledge into robot policies by fine-tuning VLMs or conditioning them with low-level control tokens. Based on their action modeling paradigms, VLA approaches can be broadly categorized into regression-based Li et al. (2023b); Brohan et al. (2023a); Liu et al. (2024c), generative/diffusion-based Black et al. (2024b); Liu et al. (2024d); Team et al. (2024), and hybrid methods Liu et al. (2025a); Ye et al. (2024b). While these unified policies show strong performance on short-horizon or single-task scenarios, they often struggle with generalization and scalability–particularly across long horizons, domains, or embodiments. To address these limitations, recent efforts have introduced hierarchical structures into end-to-end VLA models, either by modularizing perception and policy Liu et al. (2024b); Team et al. (2025b); Bjorck et al. (2025), or by structuring execution into multi-stage skills Shi et al. (2025); Physical Intelligence (2025); Raj et al. (2024). Though often labeled hierarchical, many still preserve end-to-end differentiability or latent consistency across stages. Such designs improve sample efficiency, handle long-horizon dependencies, and provide interpretable intermediate reasoning. Despite these advances, existing VLA systems remain heavily dependent on in-domain demonstrations. Some mitigate this via cross-modal or cross-embodiment data Intelligence et al. (2025); NVIDIA et al. (2025), or by scaling human video demonstrations Yang et al. (2024a); Tian et al. (2024), yet diversity and coverage remain limited. Our work enables zero-shot generalization across unseen tasks and settings, while maintaining modularity, interpretability, and execution efficiency.
@@ 7.1 Brief Summary of Methodology Advantages
Against Vanilla Video Diffusion Models-Based Methods Compared to long-horizon video diffusion models, our method operates at the primitive level with short-horizon prediction, drastically reducing computational cost and optimization difficulty. Leveraging causal distillation and self-forcing (Huang et al., 2025), we achieve real-time generation (12 FPS) in just 4 denoising steps, enabling online, on-the-fly rollout. This causal, modular design avoids error accumulation and supports dynamic, adaptive control–critical for real-world robotic deployment.
@@ B.1 The Gen6D Method
Specifically, given a generated video sequence $V=\{I_{1},I_{2},\ldots,I_{T}\}$ and a reference video $V_{r}=\{I_{r1},I_{r2},\ldots,I_{rT^{\prime}}\}$ of the robotic gripper, we estimate the gripper’s 6-DoF pose using a model-free, RGB-based pose estimator: $\mathbf{R},\mathbf{T}=PE(V,V_{r})$, where $\mathbf{R}$ and $\mathbf{T}$ denote the rotation and translation matrices, respectively. In our implementation, we use Gen6D (Liu et al., 2022b) for its simplicity and generalization capability.
@@
After generating the future rollout $x_{1:T}^{\text{img}}$ for a primitive $a_{k}$, we extract a 6-DoF end-effector trajectory $\tau_{k}=\{p_{1},\ldots,p_{T}\}$ using 6-DoF pose estimation Liu et al. (2022b) and geometric transformation. These trajectories are then mapped into Cartesian control space and executed by the robot.
@@ Appendix C Details for Hierachical Plug-and-Play Long-Horizon Compositional Generalization
We leverage LoRA-based adaptation to enable reasoning and spatial grounding, following a Chain-of-LoRA (Liu et al., 2025b)-style architecture. The structure is detailed in Figure 8.